Showing algebra....
A pile of coin 15 coins contains dimes and quaters. If the value of the pile is $1.65 find the number of dimes and quaters in the pile.

Answer by ptaylor(2198)

Let x=number of quarters
value of the quarters=0.25x
Then 15-x=number of dimes
value of the dimes=0.10(15-x)
Now we are told that the value of the quarters plus the value of the dimes equals $1.65, so:
0.25x+0.10(15-x)=1.65 get rid of parens
0.25x+1.5-0.10x=1.65 subtract 1.5 from both sides
0.25x+1.5-1.5-0.10x=1.65-1.5 collect like terms
0.15x=0.15 divide both sides by 0.15
x=1---------------------------number of quarters
15-x=15-1=14-----------------------number of dimes
CK
0.25*1+0.10*14=$1.65
0.25+1.40=1.65
1.65=1.65

Hope this helps----ptaylor